OpenStudy (jacksonjrb):

2x^(2/5)-x^(1/5)-1=0 Solve for x

OpenStudy (phi):

this is *almost* the same problem let y= x^(1/5)

OpenStudy (jacksonjrb):

So far I did y=x^(1/5) Which leaves me with 2y^2-y-1=0 and then (2y+1)(y-1)=0

OpenStudy (phi):

yes, and you get the same values for y as before

OpenStudy (jacksonjrb):

would x=y^5 ?

OpenStudy (phi):

yes, if you have \[ y= x^{\frac{1}{5}} \] raise both sides to the 5th power: \[ y^5= \left(x^{\frac{1}{5}} \right)^5 \\y^5=x^{\frac{1}{5}\cdot 5} \\ y^5= x \]

OpenStudy (phi):

notice we used the "rule" \[ \left( a^b\right)^c = a^{bc} \]

OpenStudy (jacksonjrb):

So x is equal to 1 and -0.3125

OpenStudy (phi):

(-½)^5 is not -0.3125

OpenStudy (jacksonjrb):

Are you sure? My calculator disagrees. :p

OpenStudy (jacksonjrb):

It might be an error on my part but I'm pretty sure it 's right

OpenStudy (phi):

try again.

OpenStudy (jacksonjrb):

-0.03125

OpenStudy (jacksonjrb):

My mistake

OpenStudy (phi):

yes, -1/32

OpenStudy (jacksonjrb):

I have another problem that is different then the two we did so I'll make another post. Thanks for all of your help.